# On energy-momentum conservation in non-minimal geometry-matter coupling theories

### v1

@@ Abstract
In this work, we discuss the conditions that allow the establishment of an equivalence between $f(R,T)=R+\lambda h(T)$ gravity models and General Relativity coupled to a modified matter sector. We do so by considering a $D$-dimensional spacetime and the matter sector described by nonlinear electrodynamics and/or a scalar field.
@@ I Introduction
In this paper, we aim to fill a gap in the literature by demonstrating that indeed $f(R,T)$ gravity models in which the function $f$ is given by $f(R,T)=R+\lambda h\left(T\right)$, with $h(T)$ being a function that solely depends on $T$, are physically equivalent to GR plus a modified matter sector, by considering explicit kinds of matter described by nonlinear electrodynamics (NED) and a scalar field. The paper is organized as follows. In Sec. II, we present the field equations of $f(R,T)$ gravity in their general form. In Sec. III, we perform our formal analysis by considering a $f(R,T)=R+\lambda h(T)$ gravity model coupled to NED, first particularizing $h(T)=T$ and then taking the general case. In Sec. IV, we perform the very same analysis, but this time assuming $h(T)=T$ only, and that the matter sector is solely constituted by a quintessence scalar type of matter. In Sec. V, we consider that the matter sector is composed of both NED and scalar matter. Finally, we present our main conclusions in Sec. VI.
@@ VI Summary and Conclusions
In this work, we have explored some formal aspects of the dynamics of modified theories of gravity of the $f(R,T)$ type. Even in the absence of explicit couplings between the scalars $R$ and $T$, it is generally claimed that these theories lead to violations in the conservation of the energy-momentum tensor. However, we have shown that this fact is not always so. In particular, if one considers a single matter source of scalar or electromagnetic nature in theories of the form $f(R,T)=R+\lambda h(T)$, the effect of the modified dynamics is equivalent to a modification of the matter sector, which generically becomes nonlinear. We have shown explicitly how this occurs when an arbitrary NED Lagrangian $\phi(X)$ or an arbitrary quintessence Lagrangian $F(\phi,Z)$ is taken, being the result valid in arbitrary dimension $D$. This also allowed us to see that there are some special families of theories for which the dynamics remain unchanged as compared to those found in GR. However, we also saw that when different kinds of matter are considered simultaneously, such as a NED plus a scalar sector, in nonlinear $f(R,T)$ models, there is an explicit mixing of the elements of the matter sector that prevents their interpretation as the result of a (new and nonlinear) NED plus scalar sector in GR. This is so simply because the interactions induced between the matter fields via the $h(T)$ term (of gravitational origin) cannot be re-coded in self-interactions of the participating fields. The results presented here are thus useful to better understand the nature of the mechanisms that induce new dynamics in this type of theories [28], where nonlinear self-interactions do not need to imply violations of energy-momentum conservation.

### v2

@@ Abstract
In this work, we discuss the conditions that allow the establishment of an equivalence between $f(R,T)=R+\lambda h(T)$ gravity models and General Relativity (GR) coupled to a modified matter sector. We do so by considering a $D$-dimensional spacetime and the matter sector to be described by nonlinear electrodynamics and/or a scalar field. We find that, for this particular family of models, the action and field equations can indeed be written in terms of a modified matter source within GR. However, when several matter sources are combined, this interpretation is no longer possible if $h(T)$ is a nonlinear function, due to the emergence of crossed terms that mix together the scalar and vector sectors.
@@ I Introduction
In this paper, we aim to fill a gap in the literature by demonstrating that indeed $f(R,T)$ gravity models in which the function $f$ is given by $f(R,T)=R+\lambda h\left(T\right)$, with $h(T)$ being a function that solely depends on $T$, are physically equivalent to GR plus a modified matter sector, by considering explicit kinds of matter described by nonlinear electrodynamics (NED) and a scalar field. As we will see, when the matter sector is limited to a single source, the resulting field equations are equivalent to those of GR coupled to a nonlinear version of the original matter source. However, if several matter sources are considered together and the function $h(T)$ is nonlinear on $T$, then the coupling between the scalar and vector degrees of freedom prevents a clear splitting of both sectors. In any case, it is always possible to interpret the dynamics as corresponding to GR with a conserved effective energy-momentum tensor.
The paper is organized as follows. In Sec. II, we present the field equations of $f(R,T)$ gravity in their general form. In Sec. III, we perform our formal analysis by considering a $f(R,T)=R+\lambda h(T)$ gravity model coupled to NED, first particularizing $h(T)=T$ and then taking the general case. In Sec. IV, we perform the very same analysis, but this time assuming $h(T)=T$ only, and that the matter sector is solely constituted by a quintessence scalar type of matter. In Sec. V, we consider that the matter sector is composed of both NED and scalar matter. Finally, we present our main conclusions in Sec. VI.
@@ II Field Equations of $f(R,T)$ Gravity
Note that we are working in $D$ dimensions to reinforce the generality of our analysis and results, which are by no means restricted to $4$ spacetime dimensions.
@@ VI Summary and Conclusions
In this work, we have explored some formal aspects of the dynamics of modified theories of gravity of the $f(R,T)$ type. Even in the absence of explicit couplings between the scalars $R$ and $T$, it is generally claimed that these theories lead to violations in the conservation of the energy-momentum tensor. However, we have shown that this fact is not always so. In particular, if one considers a single matter source of scalar or electromagnetic nature in theories of the form $f(R,T)=R+\lambda h(T)$, the effect of the modified dynamics is equivalent to a modification of the matter sector, which generically becomes nonlinear. We have shown explicitly how this occurs when an arbitrary NED Lagrangian $\varphi(X)$ or an arbitrary quintessence Lagrangian $F(\phi,Z)$ is taken, being the result valid in arbitrary dimension $D$. This also allowed us to see that there are some special families of theories for which the dynamics remain unchanged as compared to those found in GR. However, we also saw that when different kinds of matter are considered simultaneously, such as a NED plus a scalar sector, in nonlinear $f(R,T)=R+\lambda h(T)$ models, there is an explicit mixing of the elements of the matter sector that prevents their interpretation as the result of a (new and nonlinear) NED plus scalar sector in GR. This is so simply because the interactions induced between the matter fields via the $h(T)$ term (of gravitational origin) cannot be re-coded in self-interactions of the participating fields due to the nonlinear couplings that arise between the scalar and vector sectors of the matter terms. The results presented here are thus useful to better understand the nature of the mechanisms that induce new dynamics in this type of theories [28], where nonlinear self-interactions do not need to imply violations of energy-momentum conservation.
To conclude, one should note that for more general Lagrangians, not restricted to the decomposition $f(R,T)=R+\lambda h(T)$, the explicit non-minimal coupling between the geometric and matter sectors via products of curvature and matter, $R^{n}T^{m}$ (where $n$ and $m$ are real exponents), will in general lead to violations of energy-momentum conservation. Thus, in such a case, it is impossible to interpret the new matter terms, coming from $T^{m}$, as nonlinear realizations of the original matter action, as these terms will necessarily exchange energy and momentum with gravitons.